OpenStudy (anonymous):

what is the solution for –2(x – 3) ≥ 5 – (x + 3) ?

OpenStudy (ajprincess):

\(–2(x – 3) \ge 5 – (x + 3)\) \(-2x+6\ge5-x-3\) \(-2x+6\ge2-x\) Solve for x. Does that help? @jaderbrown

OpenStudy (anonymous):

yes, thank you. how should I graph the equation ?

OpenStudy (ajprincess):

what do u get when u solve for x?

OpenStudy (anonymous):

-4 is less than or equal to -x

OpenStudy (anonymous):

so its a closed circle and you shade it to the left or right?

OpenStudy (ajprincess):

\(-2x+6\ge2-x\) \(6\ge2-x+2x\) \(6-2\ge x\) \(4\ge x\) \(x\le4\) this means that x can take any values less than 4 and also 4. what are the values less than 4?

OpenStudy (anonymous):

3,2,1 and lower

OpenStudy (ajprincess):

in a number line these values in the right or left?

OpenStudy (anonymous):

the left, I got it now. thank you so much!

OpenStudy (ajprincess):

Welcome:)